OpenStudy (anonymous):

can u help me solve x/6 = 5/10 and how did u get you come up with the answer

OpenStudy (anonymous):

Multiply by 6. x = (5/10)*6 x = 30/10 x = 3

OpenStudy (anonymous):

could u break it down a little more im still confused thanks